OpenStudy (anonymous):

One number is 3 less than twice another. If their sum is 39, find the numbers. Which of the following systems of equations represents the word problem? a) y = 2x - 3 and y = x + 39 b) y = 2x - 3 and x + y = 39 c) y = 2(x - 3) and x + y = 39

OpenStudy (anonymous):

A MOSTLY

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line with m = 1/2 and b = 3.

OpenStudy (marissalovescats):

Say we have 2 numbers, x and y. So we know one number is 3 less then than twice another so: y= 2x-3 And we know that both numbers need to have a sum of 39 (AKA add up to 39) So x+y=39

OpenStudy (anonymous):